# Martingale deep learning for very high dimensional quasi-linear partial differential equations

### v3

@@ Abstract
High-dimensional parabolic partial differential equations (PDEs) often involve large-scale Hessian matrices, which are computationally expensive for deep learning methods relying on automatic differentiation to compute derivatives. This work aims to address this issue. In the proposed method, the PDE is reformulated into a martingale formulation, which allows the computation of loss functions to be derivative-free and parallelized in time-space domain. Then, the martingale formulation is enforced using a Galerkin method via adversarial learning techniques, which eliminate the need of computing conditional expectations in the margtingale property. This method is further extended to solve Hamilton-Jacobi-Bellman (HJB) equations and the associated Stochastic optimal control problems, enabling the simultaneous solution of the value function and optimal feedback control in a derivative-free manner. Numerical results demonstrate the effectiveness and efficiency of the proposed method, capable of solving HJB equations accurately with dimensionality up to 10,000.
@@ 1 Introduction
High-dimensional partial differential equations (PDEs) are encountered in various fields, including quantum physics, system controls, financial engineering, and data science, where traditional numerical methods become infeasable due to the curse of dimensionality (CoD) [4]. In recent years, the introduction of deep learning into numerical method designs has led to some groundbreaking advances in tackling the CoD, yielding practical approaches for high-dimensional PDEs. These approaches can be broadly classified into two categories:
@@
  • Direct approach: neural networks are trained to learn the solutions of the PDEs, as exemplified by the Physics-Informed Neural Networks (PINNs) [29; 19; 16; 33; 10], Deep Galerkin Method (DGM) [32; 1], Deep Ritz [9], Weak Adversarial Networks (WANs) [35], etc.
@@
  • Stochastic differential equation (SDE)-based approach: the PDE is reformulated into a SDE model and then solved by deep learning. Relevant works include the DeepBSDE [8; 13; 14], Forward-Backward Stochastic Neural Networks [30; 36], deep splitting method [3], deep backward schemes [21; 11], Actor-critic method [37], deep neural networks algorithms [20; 2], diffusion loss [26], diffusion Monte Carlo like approach [15], etc.
@@
When applied to high-dimensional PDEs, the above two approaches enjoy different features. In the direct approach, the empirical loss for the neural network training is generally computed over a set of randomly sampled points in the solution domain. A useful of merit of this approach is that computations for the sampled points are decoupled and thus amenable to parallel computing. However, this approach typically relies on automatic differentiation to compute the derivatives in the PDE, which can be quite expensive when the equation is truly high-dimensional and involves Hessian matrices with $d\times d$ entries. To address this issue, a stochastic dimension gradient descent (SDGD) implementation of PINNs was recently proposed to solve some very high-dimensional PDEs [19].
@@
In contrast to the direct approach, the stochastic differential equation model-based approach can avoid the expensive computation of second-order derivatives in the original PDEs by recasting the problem into a stochastic framework. However, learning stochastic models introduces new difficulties that do not appear in the direct approach: i) Most SDE-based methods, such as [8; 13; 14; 30; 36; 26; 15; 37], rely on simulating sample paths. When applied to quasilinear parabolic PDEs (see, Eq. (1)) or HJB equations (see Eq. (23)), these sample paths depend on the unknown solution or optimal control being learned, requiring iterative updates throughout the training process. These updates are sequential in time and significantly increases computational costs. ii) The SDE-based methods [3; 21; 11; 20; 2] propose to solve the problems backwardly in time, which can reduce training difficulty, but limits the time parallelizability of the algorithm.
@@
In addition to conventional high-dimensional PDEs, this work consider particularly Hamilton-Jacobi-Bellman (HJB) equations, which arise from dynamic programming methods for stochastic optimal control problems (SOCPs). The HJB equation for SOCPs contains second-order derivatives and is typically very high-dimensional, since real-world stochastic problems frequently involve numerous state and control variables. Moreover, a unique challenge in solving the HJB equation lies in the embedded minimization problem (see Eq. (23)), where the minimizer generally has no analytic expression, and the SDGD version of PINN is no longer applicable. Conventional PDE solvers need to find the minimizer for each time-space point, leading to high computational costs. To address this issue, a popular method is the policy improvement algorithm (PIA), which introduces a new neural network that serves as a feedback control to learn the minimizer; see [1] in the direct approach, and [22; 37; 20; 2] in the SDE-based approach.
@@
    Following the idea of DeepMartNet [7; 5], we introduce the system process to cast the parabolic PDE into a martingale formulation, which avoids computing all kinds of derivatives in the original problem.
@@
    The martingale formulation is carefully designed for quasilinear PDEs such that, the sample path for exploring space can be simulated in an offline manner prior to training, and the martingale property is enforced on a series of system processes simulated in parallel across time, enabling mini-batch sampling and parallel computation of the loss function across both spatial and temporal directions.
@@
    Following the approach of SOC-MartNet proposed in [6], we enforce the martingale formulation using the Galerkin method combined with adversarial learning techniques. This avoids computing the conditional expectations individually for each sampled state, and dramatically reduces the computational cost.
@@
The rest of this paper is organized as follows. In section 2, the martingale deep learning method is proposed to solve quasilinear parabolic PDEs. In section 3, the proposed method is extended to solve HJB equations. Section 4 presents the numerical results. Some concluding remarks are given in section 5.
@@ 2 Martingale deep learning for quasilinear PDEs
with a specific terminal condition $v(T,x)=g(x)$ for $x\in\mathbb{R}^{d}$. Here $\mathcal{L}$ is the differential operator as
@@
for given functions $\mu$ and $\sigma$ valued in $\mathbb{R}^{d}$ and $\mathbb{R}^{d\times q}$, respectively. The operators $\partial_{x}=\nabla_{x}$ and $\partial_{xx}^{2}=\nabla_{x}\nabla_{x}^{\top}$ are the gradient and Hessian operator, respectively, and $\operatorname{Tr}$ denotes the trace of a matrix. For a concise presentation of the main idea, the functions $\mu$, $\sigma$ and $v$ are assumed to be smooth enough to validate the involved truncation error estimates.
@@ 2.1 Martingale formulation
To present the martingale formulation of Eq. (1), we shall extend the idea of DeepMartNet proposed in [7; 5].
Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered complete probability space with $\mathbb{F}:=(\mathcal{F}_{t})_{0\leq t\leq T}$ the natural filtration of the standard $q$-dimensional Brownian motion $B:[0,T]\times\Omega\to\mathbb{R}^{q}$. We introduce a pilot process $\hat{X}:[0,T]\times\Omega\to\mathbb{R}^{d}$ used to explore space $\mathbb{R}^{d}$. A typical example of $\hat{X}$ can be given by the SDE
@@
for $t\in[0,T]$, where $\hat{\mu}(t,x):=\mu(t,x,\hat{v}(t,x))$ and $\hat{\sigma}(t,x):=\sigma(t,x,\hat{v}(t,x))$ for $(t,x)\in[0,T]\times\mathbb{R}^{d}$ with $\hat{v}$ an initial guess of $v$, and the stochastic integral with respect to $B_{s}$ is of Itô type. For $0\leq s\leq t\leq T$, let $t\mapsto X_{t}^{s}$ be the system process starting from $\hat{X}_{s}$ and generated by the operator $\mathcal{L}$, that is,
@@
We are ready to introduce the martingale formulation based on the system process $X_{t}^{s}$. For any smooth $v\in C^{1,2}$, we apply the Itô formula to $t\mapsto v(t,X_{t}^{s})$, yielding that
@@
which further implies
@@
where $R(t,x;v)$ denotes the residual error of the PDE in Eq. (1) of the function $v(t,x)$, i.e.,
@@
Under some usual conditions, the system process $\left\{X_{t}^{s}:t\in[s,T]\right\}$ is Markovian relative to the filtration $\mathbb{F}$ [31, Definition 17.2.1, Theorem 17.2.3]. Moreover, the Itô integral in Eq. (6) is an $\mathbb{F}$-martingale for $t\in[s,T]$ [27, Corollary 3.2.6], and thus it can be canceled by taking the conditional expectation $\mathbb{E}\left[\,\cdot\,|X_{s}^{s}\right]$, yielding that
@@
Letting $t$ approach $s$ and noticing that $X_{s}^{s}=\hat{X}_{s}$, we can recover the residual error $R$ from Eq. (8) by
@@
under sufficient regularity conditions, where $h>0$ is a small time step size. Eq. (7) and Eq. (9) suggest that we can solve the PDE in Eq. (1) by finding $v$ such that $\mathcal{M}_{t}^{s}$ fulfill a martingale property around $t\to s+$. Then we propose the margtingale formulation for Eq. (1):
@@
Our martingale formulation in Eq. (10) is originated from the DeepMartNet [7; 5]. In the DeepMartNet, the martingale formulation is enforced on the paths of $t\mapsto X_{t}^{0}$ given by Eq. (4) over the entire time interval $t\in[0,T]$. However, in Eq. (10), the entire diffusion $t\mapsto X_{t}^{0}$ is replaced by a series of system processes $s\mapsto X_{s}^{t}$ on the small time intervals $s\in[t,t+h]$.
This change yields the following features compared with existing works:
  • The DeepMartNet and other SDE-based approaches, such as [8; 13; 14; 30; 36; 26; 15], require the simulation of paths for $t\mapsto X_{t}^{0}$. This involves solving Eq. (4) with $s=0$ over the entire time interval $t\in[0,T]$. Moreover, Eq. (4) relies on the unknown function $v$ under approximation, which means that Eq. (4) needs to be re-computed whenever the approximation of $v$ is updated. This re-computing process is sequential in time and imposes substantial computational costs.
  • For the martingale formulation in Eq. (10), the simulations of the system processes $t+h\mapsto X_{t+h}^{t}$ are decoupled across the time steps $t\in[0,T-h]$. This enables mini-batch sampling of the time steps $t\in[0,T-h]$, allowing the martingale property in Eq. (10) to be enforced individually and in parallel for each sampled time step. This feature significantly speed up the computation of our method.
@@ 2.2 Martingale formulation combined with a Galerkin method
Compared with the original PDE in Eq. (1), the desirable feature of Eq. (10) is derivative-free, which constitute a fundamental improvement over the DeepMartNet and SOC-MartNet proposed in [7; 5; 6]. However, this feature comes at the cost of approximating the conditional expectation for samples of $\hat{X}_{t}$. To address this issue, we shall reformulate Eq. (10) into a weak form.
Let $\mathcal{T}$ be the set of test functions consisting of all bounded and smooth function on $[0,T]\times\mathbb{R}^{d}$, then Eq. (10) can be ensured by
$\int_{0}^{T-h}\mathbb{E}\big{[}\rho(t,\hat{X}_{t})\mathbb{E}\left[\mathcal{M}_{t+h}^{t}|\hat{X}_{t}\right]\big{]}\,\mathrm{d}t=0$ (11)
for any $\rho\in\mathcal{T}$. Since $\rho(t,\hat{X}_{t})$ is $\hat{X}_{t}$-measurable, it can be moved into $\mathbb{E}\left[\;\cdot\;|\hat{X}_{t}\right]$, yielding that
@@
The approximation error of Eq. (12) is of order $O(h^{2})$, following from the facts that the Euler scheme applied to Eq. (4) enjoys a local truncation error of weak-2nd order [23, Propostion 5.11.1] and the left-rectangular quadrature formula applied to the integral in Eq. (5) is of 2nd order, too.
@@ 2.3 Adversarial learning for the martingale formulation
The minimax problem in Eq. (14) can be naturally solved by adversarial learning as in [35].
Specifically, we approximate the functions $v$ and $\rho$ by neural networks $v_{\theta}$ and $\rho_{\eta}$ parameterized by $\theta$ and $\eta$, respectively. To fulfill the terminal condition $v(T,x)=g(x)$ for $x\in\mathbb{R}^{d}$, the network $v_{\theta}$ takes the form of
@@
The adversarial network $\rho_{\eta}$ plays the role of test functions. By our experiment results, $\rho_{\eta}$ is not necessarily to be very deep, but instead, it can be a shallow network with enough output dimensionality. Following the multiscale neural network approach in [25; 24], we consider a typical $\rho_{\eta}$ in the form of
@@
The neural networks can be trained to fit Eq. (14) by stochastic gradient algorithms with the expectation in Eq. (15) approximated by mini-batch sampling. Specially, we introduce a uniform time partition on $[0,T]$, i.e., $t_{n}=nh$ for $n=0,1,\cdots,N$ with the step size $h=T/N$. By applying the Euler scheme to Eq. (3), the sample paths of the pilot process $\hat{X}$ are generated by
@@
for $n=0,1,\cdots,N-1$ and $m=1,2,\cdots,M$, where all $\xi_{n}^{m}$ are i.i.d. samples from $\mathrm{N}(0,I_{q})$. Then we define the empirical version of $G$ as
@@
where $\mathcal{M}$ is given in Eq. (13), and $A$ is a index subset randomly taken from $\{0,1,\cdots,N-1\}\times\{1,2,\cdots,M\}$. Further, the loss function in Eq. (14) can be estimated by
@@
Here, inspired by [12; 19], the sets $A_{1}$ and $A_{2}$ are disjoint to ensure the mini-batch estimation in Eq. (21) is unbiased.
@@
  • Derivative-free approach: Different from the deep learning methods directly applied on Eq. (1), Algorithm 1 is free of computing any derivatives by automatic differentiation, espeically the $d\times d$-dimensional Hessian matrix $\partial_{xx}^{2}v$, which generate much gain in efficiency for problems with very high spatial dimensionality.
  • Parallel loss computation: The sample paths $\left\{\hat{X}_{n}^{m}\right\}_{n=0}^{N}$ are independent of any unknown quantity to be learned, and can be simulated offline using Eq. (19) prior to training. Throughout training, all the summation terms in Eq. (20) can be computed parallelly even in the time direction. This feature differs significantly from existing SDE model-based deep learning methods, whose sample paths typically need to be updated throughout training, in a sequential fashion over time.

### v4

@@ Abstract
In this paper, a highly parallel and derivative-free martingale neural network learning method is proposed to solve Hamilton-Jacobi-Bellman (HJB) equations arising from stochastic optimal control problems (SOCPs), as well as general quasilinear parabolic partial differential equations (PDEs). In both cases, the PDEs are reformulated into a martingale formulation such that loss functions will not require the computation of the gradient or Hessian matrix of the PDE solution, while its implementation can be parallelized in both time and spatial domains. Moreover, the martingale conditions for the PDEs are enforced using a Galerkin method in conjunction with adversarial learning techniques, eliminating the need for direct computation of the conditional expectations associated with the martingale property. For SOCPs, a derivative-free implementation of the maximum principle for optimal controls is also introduced. The numerical results demonstrate the effectiveness and efficiency of the proposed method, which is capable of solving HJB and quasilinear parabolic PDEs accurately in dimensions as high as 10,000.
@@ 1 Introduction
High-dimensional partial differential equations (PDEs) are encountered in various fields, including quantum physics, system controls, financial engineering, and data science, where conventional approaches [37, 30, 34, 3, 6, 38, 44, 46, 45, 40] suffer from the curse of dimensionality (CoD) [5], that is, the computation complexity increases exponentially with the dimension of the problems. In recent years, the introduction of deep learning into numerical method designs has led to some groundbreaking advances in tackling the CoD, yielding practical approaches for solving high-dimensional PDEs. These approaches can be broadly classified into two categories:
@@
  • Direct approach: neural networks are trained to learn the solutions of the PDEs, as exemplified by the Physics-Informed Neural Networks (PINNs) [32, 21, 18, 39, 12], Deep Galerkin Method (DGM) [36, 1], Deep Ritz [11], Weak Adversarial Networks (WANs) [42], etc.
@@
  • Stochastic differential equation (SDE)-based approach: the PDE is reformulated into a SDE model and then solved by deep learning. Relevant works include DeepBSDE [10, 15, 16], forward-backward stochastic neural networks [33, 43], deep splitting method [4], deep backward schemes [23, 13], actor-critic method [47], deep neural networks algorithms [22, 2], diffusion loss [28], diffusion Monte Carlo-like approach [17], etc.
@@
When applied to high-dimensional PDEs, the above two approaches enjoy different features. In the direct approach, the empirical loss for the neural network training is generally computed over a set of randomly sampled points in the solution domain. A useful merit of this approach is that computations for the sampled points are decoupled and thus amenable to parallel computing. However, this approach typically relies on automatic differentiation to compute the derivatives in the PDE, which can be quite expensive when the PDE is truly high-dimensional and involves Hessian matrices with $d\times d$ entries. To address this issue, a stochastic dimension gradient descent (SDGD) implementation of the PINNs was recently proposed to solve some very high-dimensional PDEs [21].
@@
In contrast to the direct approach, the SDE model-based approach can avoid the expensive computation of second-order derivatives in the original PDEs by recasting the problem into a stochastic framework. However, learning stochastic models introduces new difficulties that do not appear in the direct approach: i) Most SDE-based methods, such as [10, 15, 16, 33, 43, 28, 17, 47], rely on simulating sample paths. When applied to quasilinear parabolic PDEs (see, Eq. (1)) or HJB equations (see Eq. (23)), these sample paths depend on the unknown solution or optimal control being learned, requiring iterative updates throughout the training process. These updates are sequential in time and significantly increase computational costs. ii) The SDE-based methods in [4, 23, 13, 22, 2] propose to solve the problems backwardly in time, which can reduce training difficulty, but limits the time parallelizability of the algorithm.
@@
In addition to conventional high-dimensional PDEs, this work considers particularly Hamilton-Jacobi-Bellman (HJB) equations, which arise from dynamic programming methods for stochastic optimal control problems (SOCPs). The HJB equation for SOCPs contains second-order derivatives and can be very high-dimensional as real-world stochastic problems frequently involve numerous state and control variables. Moreover, a unique challenge in solving the HJB equation lies in the embedded minimization problem (see Eq. (23)), where the minimizer generally has no analytic expression and the SDGD version of PINN [21] is no longer applicable. Taditional PDE solvers need to compute the minimizer for each time-space point, leading to high computational costs. To address this issue, a popular method is the policy improvement algorithm (PIA), which introduces a new neural network for a feedback control to learn the minimizer; see [1] in the direct approach, and [24, 47, 22, 2] in the SDE-based approach.
@@
    Following the idea of DeepMartNet [9, 7], we introduce a system process to cast the parabolic PDE into a martingale formulation, which avoids computing all kinds of derivatives in the original problem.
@@
    The martingale formulation is carefully designed for quasilinear PDEs such that, the sample path for exploring space can be simulated in an offline manner prior to training, and the martingale property is enforced on a series of system processes simulated in parallel at different time instances, enabling mini-batch sampling and parallel computation of the loss function across both spatial and temporal directions.
@@
    Following the approach of SOC-MartNet proposed in [8], we enforce the martingale formulation using a Galerkin method combined with adversarial learning techniques. This avoids computing the conditional expectations individually for each sampled state, dramatically reduces computational cost, and improves the robustness of the algorithm.
@@
The rest of this paper is organized as follows. In section 2, the martingale deep learning method is proposed to solve quasilinear parabolic PDEs. In section 3, the proposed method is extended to solve HJB equations arising from SOCPs. Section 4 presents the numerical results. Some concluding remarks are given in section 5.
@@ 2 Martingale deep learning for quasilinear PDEs
with a specific terminal condition $v(T,x)=g(x)$ for $x\in\mathbb{R}^{d}$. Here $\mathcal{L}$ is the differential operator defined as
@@
for some given functions $\mu$ and $\sigma$ valued in $\mathbb{R}^{d}$ and $\mathbb{R}^{d\times q}$, respectively. The operators $\partial_{x}=\nabla_{x}$ and $\partial_{xx}^{2}=\nabla_{x}\nabla_{x}^{\top}$ denote the gradient and Hessian operator, respectively, and $\operatorname{Tr}$ denotes the trace of a matrix. For a concise presentation of the main idea, the functions $\mu$, $\sigma$ and $v$ are assumed to be smooth enough to validate the involved truncation error estimates.
@@ 2.1 Martingale formulation
To present the martingale formulation of Eq. (1), we shall extend the idea of DeepMartNet proposed in [9, 7]. Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered complete probability space with $\mathbb{F}:=(\mathcal{F}_{t})_{0\leq t\leq T}$ being the natural filtration of the standard $q$-dimensional Brownian motion $B:[0,T]\times\Omega\to\mathbb{R}^{q}$. We introduce a pilot process $\hat{X}:[0,T]\times\Omega\to\mathbb{R}^{d}$ to be used to explore space $\mathbb{R}^{d}$. A typical example of $\hat{X}$ can be given by the SDE
@@
for $t\in[0,T]$, where $\hat{\mu}(t,x):=\mu(t,x,\hat{v}(t,x))$ and $\hat{\sigma}(t,x):=\sigma(t,x,\hat{v}(t,x))$ for $(t,x)\in[0,T]\times\mathbb{R}^{d}$ with $\hat{v}$ as an initial guess of $v$, and the stochastic integral with respect to $B_{s}$ is of the Itô type. For $0\leq s\leq t\leq T$, let $t\mapsto X_{t}^{s}$ be the system process starting from $\hat{X}_{s}$ (the state of the pilot process at time $s$) and generated by the operator $\mathcal{L}$, that is,
@@
We are ready to introduce the martingale formulation based on the system process $X_{t}^{s}$. For any smooth $v\in C^{1,2}$, we apply the Itô formula to $t\mapsto v(t,X_{t}^{s})$, yielding
@@
which leads to the following process
@@
where $R(t,x;v)$ denotes the residual error of the PDE in Eq. (1) for the function $v(t,x)$, i.e.,
@@
Under some usual conditions, the system process $\left\{X_{t}^{s}:t\in[s,T]\right\}$ is Markovian relative to the filtration $\mathbb{F}$ [35, Definition 17.2.1, Theorem 17.2.3]. Moreover, the Itô integral in Eq. (6) is an $\mathbb{F}$-martingale for $t\in[s,T]$ [29, Corollary 3.2.6], and thus it can be eliminated by taking the conditional expectation $\mathbb{E}\left[\,\cdot\,|X_{s}^{s}\right]$, yielding that
@@
Letting $t=s+h$ and noticing that $X_{s}^{s}=\hat{X}_{s}$, we can recover the residual error $R$ from Eq. (8) by
@@
under sufficient regularity conditions, where $h>0$ is a small time step size. Eq. (7) and Eq. (9) suggest that we can solve the PDE in Eq. (1) by finding $v$ such that $\mathcal{M}_{t}^{s}$ fulfill a martingale property around $t\to s+$. Then, we can propose the margtingale formulation for Eq. (1):
@@
The martingale formulation in Eq. (10) originates from our previous work on DeepMartNet [9, 7]. In the DeepMartNet, the martingale formulation is enforced on the paths of $t\mapsto X_{t}^{0}$ given by Eq. (4) over the entire time interval $t\in[0,T]$. However, in Eq. (10), the entire diffusion $t\mapsto X_{t}^{0}$ is replaced by a series of system processes $s\mapsto X_{s}^{t}$ on small time intervals $s\in[t,t+h]$.
This change yields the following important features compared with the previous works:
  • The DeepMartNet and other SDE-based approaches, such as [10, 15, 16, 33, 43, 28, 17], require the simulation of paths for $t\mapsto X_{t}^{0}$. This involves solving Eq. (4) with $s=0$ over the entire time interval $t\in[0,T]$. Moreover, Eq. (4) relies on the unknown function $v$ under approximation, which means that Eq. (4) needs to be re-computed whenever the approximation of $v$ is updated. This re-computing process is sequential in time and imposes substantial computational costs due to the lack of parallelizablity.
  • For the martingale formulation in Eq. (10), the simulations of the system processes $t+h\mapsto X_{t+h}^{t}$ are decoupled across the time steps $t\in[0,T-h]$. This enables mini-batch sampling of the time steps $t\in[0,T-h]$, allowing the martingale property in Eq. (10) to be enforced individually and in parallel for each sampled time step. This feature significantly speed up the computation of the method.
@@ 2.2 Martingale formulation combined with a Galerkin method
Compared with the original PDE in Eq. (1), the desirable feature of Eq. (10) is derivative-free, which constitutes a fundamental improvement over the DeepMartNet and SOC-MartNet proposed in [9, 7, 8]. However, the approach still requires approximating the conditional expectation for samples of ${X}_{t}^{s}$. To address this issue, we shall reformulate Eq. (10) into a weak form.
Let $\mathcal{T}$ be the set of test functions consisting of all bounded and smooth function on $[0,T]\times\mathbb{R}^{d}$, then Eq. (10) can be ensured by requiring
$\int_{0}^{T-h}\mathbb{E}\left[\rho(t,\hat{X}_{t})\mathbb{E}\left[\mathcal{M}_{t+h}^{t}|\hat{X}_{t}\right]\right]\,\mathrm{d}t=0$ (11)
for any $\rho\in\mathcal{T}$. Since $\rho(t,\hat{X}_{t})$ is $\hat{X}_{t}$-measurable, it can be moved into $\mathbb{E}\left[\;\cdot\;|\hat{X}_{t}\right]$, yielding
@@
The approximation error of Eq. (12) is of order $O(h^{2})$, following from the facts that the Euler-Maruyama scheme applied to Eq. (4) enjoys a local truncation error of weak-2nd order [25, Propostion 5.11.1] and the left-rectangular quadrature formula applied to the integral in Eq. (5) is of 2nd order as well.
@@ 2.3 Adversarial learning for the martingale formulation
The minimax problem in Eq. (14) can be naturally solved by adversarial learning as in [42]. Specifically, we approximate the functions $v$ and $\rho$ by neural networks $v_{\theta}$ and $\rho_{\eta}$ parameterized by $\theta$ and $\eta$, respectively. To fulfill the terminal condition $v(T,x)=g(x)$ for $x\in\mathbb{R}^{d}$, the network $v_{\theta}$ takes the form of
@@
The adversarial network $\rho_{\eta}$ plays the role of test functions in classical Galerkin methods for solving PDEs. By our experiment results, $\rho_{\eta}$ is not necessarily to be very deep, but instead, it can be a shallow network with enough output dimensionality. Following the multiscale neural network approach in [27, 26], we consider a typical $\rho_{\eta}$ in the form of
@@
The neural networks can be trained to fit Eq. (14) by stochastic gradient algorithms with the expectation in Eq. (15) approximated by a mini-batch sampling. Specially, we introduce a uniform time partition on $[0,T]$, i.e., $t_{n}=nh$ for $n=0,1,\cdots,N$ with the step size $h=T/N$. By applying the Euler scheme to Eq. (3), the sample paths of the pilot process $\hat{X}$ are generated by
@@
for $n=0,1,\cdots,N-1$ and $m=1,2,\cdots,M$, where all $\xi_{n}^{m}$ are i.i.d. samples from $\mathrm{N}(0,I_{q})$. Then, we define the empirical version of $G$ as
@@
where $\mathcal{M}$ is given in Eq. (13), and $A$ is a index subset randomly taken, wthout replacements, from $\{0,1,\cdots,N-1\}\times\{1,2,\cdots,M\}$. Further, the loss function in Eq. (14) can be estimated by
@@
Here, inspired by [14, 21], the sets $A_{1}$ and $A_{2}$ are disjoint to ensure the mini-batch estimation in Eq. (21) is unbiased.
@@
  • Derivative-free approach: Different from the deep learning methods directly applied on Eq. (1), Algorithm 1 is free of computing any derivatives by automatic differentiation, espeically the $d\times d$-dimensional Hessian matrix $\partial_{xx}^{2}v$, and gains much efficiency for problems with very high spatial dimensionality.
  • Parallel loss computation: The sampled pilot paths $\{\hat{X}_{n}^{m}\}_{n=0}^{N}$ are independent of any unknown quantity to be learned, and can be simulated offline using Eq. (19) prior to training. Throughout training, all the summation terms in Eq. (20) can be computed parallelly even in the time direction. This feature differs significantly from existing SDE model-based deep learning methods, whose sample paths typically need to be updated throughout training, in a sequential fashion over time direction.